A child uses a vibrator S that hits the surface of the water in a pond at a rate of 50 vibrations in 5s. An angler is found not far from the vibrator.

Determine the frequency and the period of the vibrator


1
Expert's answer
2021-03-23T11:10:51-0400

By definition, the frequency is the number of vibrations in one second. Thus, since it makes 50 vibrations in 5s, in one second it will make:


"f = \dfrac{50}{5s} = 10\dfrac{1}{s} = 10Hz"

By defintion, the period is:


"T = \dfrac{1}{f} = \dfrac{1}{10Hz} = 0.1s"

Answer. Frequency: 10Hz, Period: 0.1s.